OpenStudy (anonymous):

what is th final speed of an object n how far does it travel accelerating if its initial speed is 10m/s n it accelerates at 8m/s2 for 5s???? Simple question CAN any one

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

help

mathslover (mathslover):

\[\huge{v=u+at}\] \[\huge{v=10ms^{-1}+8ms^{-2}({5s})}\] \[\huge{v=10ms^{-1}+40ms^{-1}}\] \[\huge{v=50ms^{-1}}\]

mathslover (mathslover):

can u calculate s ( displacement ) by the following formula : \[\huge{s=ut+\frac{1}{2} (a)t^2}\]

mathslover (mathslover):

just put all values that u have been given : \[\huge{s = 50 m + 100m }\] \[\huge{s=150m}\]

OpenStudy (anonymous):

thankyou lover

OpenStudy (anonymous):

u r a gr888 helper

mathslover (mathslover):

my pleasure @Master.RohanChakraborty